Suppose {p} is a cauchy sequence in a metric space X and some subsequence {p} converges to a point a belongs to X . Prove that the full sequence {p} converges to a .
1
Expert's answer
2019-07-05T11:01:33-0400

Suppose "\{{p_n} \}" is a cauchy sequence in a metric space X and some subsequence "\{p_{n_k}\}" converges to a point "a" belongs to "X".

Let "\epsilon > 0." There exists "N" such that, for "n, m \geq N , d(p_n, p_m) <\epsilon" ( "\{{p_n} \}" is a cauchy sequence ). Then, consider any "n \geq N" :

for "k" sufficiently large (so that "n_k \geq N," ), "d(p_n, p_{n_k} ) \leq \epsilon" .

Taking the limit as "k \rightarrow \infty", it follows that "d(p_n, a)\leq \epsilon" .

Or: if "k" is sufficiently large then "d(p_{n_k}, a) <\epsilon" (by the assumption "p_{n_k }\rightarrow a" ), so "d(p_n, a) \leq d(p_n, p_{n_k}) + d(p_{n_k}, a) < 2 \epsilon".

In any case, we conclude that "d(p_n, a)" becomes arbitrarily small for "n" large, i.e. "p_n \rightarrow a" .